# DeepInsert: Early Layer Bypass for Efficient and Performant Multimodal Understanding

### v2

@@ Abstract
The hyperscaling of data and parameter count in transformer models is yielding diminishing performance improvement, especially when weighed against training costs. Such plateauing underlines a growing need for more efficient finetuning and inference, without sacrificing performance. This is particularly pressing for multimodal learning, where the overhead of processing multimodal tokens alongside language data often limits the practical viability of these systems. In parallel, advances in representation learning and interpretability have deepened our understanding of how such models process and encode information. Notably, recent work has uncovered implicit cross-modal alignment in the deeper layers of large pretrained models. Interestingly, this aligns with our own observations that models naturally defer most cross-modal token interactions to deeper stages of computation. Building on this, we propose a simple modification. Instead of concatenation with the language prompt at the start, we insert multimodal tokens directly into the middle, allowing them to entirely bypass the early layers. Our results with diverse modalities: 1) LLaVA & BLIP for vision, 2) LTU for audio, and 3) MoLCA for molecular data, indicate that our method reduces computational costs during both training and inference, while at the very least, preserving, if not surpassing the performance of existing baselines. Our work has important implications for scaling and composing pretrained models in a resource-efficient manner.
@@ 1 Introduction
Advances in large language models (LLMs) [6, 13, 79, 63] have spurred the development of general purpose reasoning models capable of multimodal comprehension in diverse application domains. Although scaling has been crucial for this success [39, 87, 32], the growing awareness of the enormous costs associated with large multimodal language models (MLMs/MLLMs) [74, 30] has placed focus on methods for efficient finetuning of pretrained models and speeding up their inference [33, 43]. For multimodal tasks such as vision-language understanding [67, 44, 50], the use of pretrained unimodal models, rather than training from scratch, has proven effective at reducing training overhead [66]. However, overall efficiency remains a challenge, as multimodal inputs impose additional computational costs during both training and inference, particularly when processed alongside language prompts.
Alongside scaling, a major key to general purpose multimodal training is effective cross-modal alignment. In practice, this is usually achieved with a combination of pretraining stages with paired multimodal data, followed by multitask instruction finetuning for specialized downstream tasks [15]. Preference tuning techniques, such as reinforcement learning from human feedback (RLHF) [64] can produce further improvements [76, 81]. However, most works in this area tend to propose modifications to the training algorithm and/or choice of dataset, while largely treating the pretrained components as replaceable black-box models. To complement this body of work, we begin our investigation with a bottom-up approach by analyzing what happens inside these models. For this qualitative analysis, we rely on pretrained VLMs as a starting point.
Within the literature on analyzing representations and interpretability, our initial inspiration comes from Huh et al. [36], who argue that representations from different unimodal models converge towards a universal platonic representation. Although the platonic hypothesis itself is contentious, the underlying evidence suggests that model representations across modalities show increasing alignment in the deeper layers (a quick glance at Figs. 4, 5 in the Appendix reaffirms it). This, coupled with recalling the inner product mechanism of self-attention, instantly invites the next question: in a multimodal setting, are cross-modal tokens likely to interact more in the deeper layers? Analyzing the attention maps suggests that the bulk of the activity is, indeed, confined to the middle and latter layers (see Fig. 2 and results in concurrent work [38]). All of the preceding exposition thus culminates in our central question: Are the early layers of the LLM really necessary for multimodal token processing, and if not, can we remove redundancy without sacrificing performance? Our key contributions can then be summarized as follows.
  1. 1.
    We show that letting multimodal tokens bypass early layers significantly reduces training and inference compute, with virtually no loss in performance—and in some cases, even marginal gains—in an almost free lunch fashion. We refer to this framework as DeepInsert.
  2. 2.
    We comprehensively demonstrate DeepInsert’s utility across modalities in achieving the desired goals, via replicable experiments on popular open-source MLLMs: vision (LLaVA, BLIP [49, 45]), audio (LTU [28]), and molecules (MolCA [52]). We also reveal a natural performance-efficiency tradeoff (see Fig. 1), enabling practitioners to select models based on their specific needs.
@@
The remainder of the paper is organized as follows. Sec. 2 begins by placing our work in the context of broader literature. Sec. 3 motivates our framework, while Sec. 4 details the general notion of the architectural modification, highlighting implementation and design considerations. Experiments in Sec. 5 demonstrate the versatility and effectiveness across a variety of popular open source multimodal models. Sec. 6 concludes with promising directions for broader application. In addition to our more practical contributions, our work hints at novel underlying mechanisms for multimodal processing, which may be of independent interest.
@@ 2 Related Work
MLLMs & Alignment: The general architecture of MLLMs considered in our work comprises a pretrained LLM, frozen perceptual encoders for different modalities (image, audio, video, etc. [26, 67, 75, 90]), and a trainable mapping module (can be a lightweight MLP or a larger Perceiver [37]) that aligns the modality with the LLM input space. The LLM prompts consist of text and multimodal tokens, the latter of which are obtained from the encoder via the mapping module. This general regime encompasses a wide range of works in all modalities [1, 45, 88, 3, 49, 28, 25, 60], and thus, our architectural modification is fairly universal in terms of applicability.
@@
Based on where alignment is achieved, our method relates best to the idea of model stitching [42, 61], that is, grafting initial layers of a neural network onto the latter half of another with an affine layer. Our work can thus be interpreted as multimodal model stitching for efficiency, without any additional bells and whistles.
@@
Efficiency & Layer Skipping: Low-rank adapters (LoRA) [34, 16] are commonly used for computationally-efficient finetuning of LLM-based models. The low-rank constraint enables finetuning large models with only a small set of additive parameters, reducing memory usage. Note, however, that inference remains just as costly. An alternative is model compression; an existing model can be effectively distilled, leading to accelerated inference and almost-at-par capabilities [22, 80]. But while these approaches exploit inherent parametric redundancies in LLMs, our approach, as we shall demonstrate, is motivated via functional redundancy and, as such, largely complementary.
The other line of work pertains to layer skipping, which has been employed in LLMs for speculative decoding and speeding up inference [14, 21, 18]. More recently, its application to MLLMs has been considered in [71, 85]. Specifically, Shukor and Cord [71] combine layer skipping with compression and pruning of the base LLMs and demonstrate impressive efficiency gains in their custom set-up. However, it is unclear how well their method scales as their models, when trained in the LLaVA setup, are unable to match baseline performance. On the other hand, concurrent work by Zeng et al. [85] proposes adaptive token skipping to speed up training and inference. However, their token selection is ad hoc and LLaVA-centric. It is unclear if it generalizes to other VLMs or other modalities.
@@
Layer-Skipping Approaches. A number of recent works have proposed layer-skipping or depth-adaptive strategies to improve inference efficiency in large models. Methods such as FlexiDepth [54] and AdaSkip [56] dynamically adjust the number of transformer layers executed based on input complexity, while MoLe-VLA [89] and $\gamma$-MoD [55] introduce mixture-of-expert or modality-specific modules to selectively bypass computation. These approaches, while effective within their respective domains, are often tailored to unimodal LLMs (e.g., FlexiDepth [54], AdaSkip [56]) or require additional architectural modifications for multimodal settings (e.g., MoLe-VLA [89], $\gamma$-MoD [55]). In contrast, DeepInsert does not rely on context-dependent routing or auxiliary modules. Instead, it provides a principled framework for inserting multimodal features at deeper layers, enabling full layer-skipping of the initial stages without degrading performance. Importantly, DeepInsert maintains strong performance across a wide spectrum of modalities—including vision, audio, and molecular data—while preserving the same training setup.
@@
Token-Reduction Approaches. Another line of research focuses on reducing the number of input tokens processed during inference. Representative examples include ST3 [93], Skip-Vision [86], FastV [10], VTW [48], PruMerge [69], MADTP [7], and PACT [17], which prune redundant tokens or learn adaptive token selection strategies. These methods are largely motivated by the redundancy of visual tokens in vision–language models. DeepInsert is conceptually orthogonal to this direction: token-reduction can be applied on top of DeepInsert to further improve efficiency. For instance, we implemented FastV [10] and VTW [48] with DeepInsert and observed that DeepInsert+FastV/VTW maintains performance comparable to or better than baseline+FastV/VTW. This suggests that DeepInsert is particularly robust when combined with token pruning, and can yield additional efficiency gains beyond baseline models. Moreover, training-free methods such as PruMerge [69] also exhibit lower performance declines when applied to DeepInsert models compared to standard LLaVA, reinforcing the complementarity of these two paradigms.
@@ 3 Where Multimodal Tokens Matter
This section motivates our proposed framework. As given in our introduction, we first explore the nature of representational alignment in pretrained unimodal models, in the setup of Huh et al. [36]. However, it remains difficult to quantify exactly how this alignment could be beneficial and thus for brevity, we relegate it to Appendix A.1. For concrete evidence, we propose analyzing the attention maps of pretrained MLLMs. The inspiration comes from works on mechanistic interpretability, which suggests that the forward pass of an LLM can be broadly delineated into phases: a) the function determination phase that processes the prompt to determine the task that needs to be executed (for instance, get_capital()) and b) the execution of the function for the given input query (for instance, get_capital(France) = Paris) [31, 57, 59]. These observations lead us to ask whether a similar delineation holds for LLM-based MLLMs: specifically, if we interpret the multimodal tokens as the input query, can we identify where the analogous query execution takes place? Note that query execution would take an inherently different form, since multimodal tokens are generally either prepended, or inserted in the middle of language prompt. Therefore, attention maps are a natural proxy here for understanding how visual tokens move information to the language token for solving multimodal tasks. For this study, we use the LLaVA v1.5 model [49] as our testbed.
Our approach is as follows. We start with a balanced subset of multiple choice type questions (MCQs) drawn from the Visual7W dataset [92], which draws images from COCO [11]. Using MCQ-based questions allows us to first isolate data samples where the model succeeds in answering correctly by comparing the token logit against the choices (A, B, C, or D) and then, obtain the relevant attention activity for these examples, both in just one forward pass. Note that isolating the correct instances is important, since instances of hallucination can significantly alter the attention behavior [38]. The corresponding attention scores from the visual token to the last language token responsible for prediction, as a function of model depth, can then be visualized either by averaging across attention heads, or isolating average behavior for the top-$k$ attention heads per layer. Finally, note that we are interested in layer-relative contribution per visual token. Therefore, we normalize the contribution in each layer by its net contribution across layers. This allows us to cleanly visualize both tokens that contribute high absolute value of attention scores (such as CLIP summary tokens [62]) with those that do not. Finally, for ease of visualization, we randomly subsample image tokens (100 out of 576 tokens) and arrange them left to right in order of increasing positional indices. In the main exposition, we visualize attention with top-$k$ ($k=5$ heads for each layer), in Fig. 2. and relegate the other cases (aggregate over heads, with/without normalization) along with some secondary details, to Appendix A.2.
It is evident from the figure that that vision tokens tend to interact with language predominantly in the intermediate layers. Furthermore, the notion seems to hold for other visualization choices (see Appendix A.2), lending credence to our intuition that vision-to-language information transfer is stronger in intermediate layers and that the earlier layers are somewhat redundant when it comes to performing this effectively. We note that a concurrent line of experimentation reveals similar findings [38], albeit with the motivation to study object hallucination.
@@
Figure 2: Illustration of vision to language token activation in LLaVA-7B. The y-axis represents LLM depth, the x-axis consists of vision tokens arranged in increasing order of positional indices. The score represents the relative layerwise contribution for that token, based on the top-5 attention heads per layer. The figure suggests that the latter layers contain the majority of attention activity.
@@
Remark. In the complete layer-wise depiction (Fig. 6 in the Appendix), there is a strong activation pattern in the first two layers that dominates all other activity. However, we believe that this interaction is unlikely to be the key point of vision-to-language information transfer and we offer two key pieces of evidence for this. Firstly, unlike activation in the middle layers, [4, 38] suggest that this behavior is model-specific (i.e., LLaVA), and is not consistently observed for other choices of VLMs/encoders. Secondly, our own experiments in the Appendix B suggest that even in the base LLaVA model with no pretraining, these layers can be skipped without significant drops in performance (at least to the extent suggested by the activity). Consequently, we do not include the first two layers in Fig. 2 to keep the visualization cleaner.
@@ 4 Proposed Framework
Our observations above bring us back to the central question of our study: Do we really need the early LLM layers to parse vision (or more generally, multimodal) information? That is, can we limit multimodal processing to only later layers, without sacrificing performance? To investigate, we propose our simple yet effective modification for general MLLM frameworks: inserting the multimodal tokens directly into the intermediate layers of the LLM. In some sense, one can view it as finding an alternative alignment in the models’ respective semantic representation spaces, akin to model stitching. We call this DeepInsert and illustrate the general approach in Fig. 3. Before jumping into experiments, we first highlight (in our opinion) some important considerations for readers looking to implement this in practice.
(a) Conventional architecture for popular MLLMs.

### v3

@@ Abstract
Hyperscaling of data and parameter count in LLMs is yielding diminishing improvement when weighed against training costs, underlining a growing need for more efficient finetuning and inference without sacrificing performance. This is especially so for multimodal language models (MLMs), where the overhead of processing multimodal tokens can limit their practical viability. Parallely, recent work has uncovered implicit cross-modal alignment in the deeper layers of large MLMs, deepening our understanding of how MLMs process and encode information. Motivated by this, and our observation that MLMs naturally defer most cross-modal token interactions to deeper layers of the model, we propose a simple modification. Instead of concatenation with the language prompt at the start, we insert multimodal tokens directly into the middle, allowing them to entirely bypass the early layers. Our results with diverse modalities, (i) LLaVA & BLIP for vision, (ii) LTU for audio, and (iii) MoLCA for molecular data, and model sizes, starting from 350M to 13B parameters, indicate that our method reduces both training and inference costs, while at least preserving, if not surpassing the performance of existing baselines.
DeepInsert: Early Layer Bypass for Efficient and Performant Multimodal Understanding
Moulik Choraria†,∥,thanks: Correspondence: moulikc2@illinois.edu; Equal contribution; Code: https://github.com/MoulikChoraria/DeepInsert          Xinbo Wu†,∥          Akhil Bhimaraju†,∥
Nitesh Sekhar§          Yue WuΦ          Xu Zhang§          Prateek Singhalψ          Lav R. Varshneyζ
UIUC, §Amazon, ΦCapital One, ψApple, ζ Stony Brook University
@@ 1 Introduction
Advances in large language models (LLMs) (Brown et al., 2020; Chung et al., 2022; Touvron et al., 2023; OpenAI et al., 2024) have spurred the development of models capable of multimodal comprehension in diverse application domains. While scaling has been crucial for this success (Kaplan et al., 2020; Zhai et al., 2021; Hoffmann et al., 2022), the computational costs associated with large multimodal language models (MLMs/MLLMs) (Strubell et al., 2019; Grattafiori et al., 2024) necessitates methods for efficient finetuning and inference (Hu et al., 2021; Leviathan et al., 2023). For multimodal tasks such as vision-language understanding (Radford et al., 2021; Li et al., 2022; Liu et al., 2024c), the use of pretrained unimodal models, rather than training from scratch, is effective at reducing training overhead (Qi et al., 2024). However, overall efficiency remains a challenge, as multimodal inputs impose additional computational costs, when processed alongside language prompts.
Another key to multimodal training is effective cross-modal alignment, which is achieved with a combination of pretraining stages, followed by multitask instruction finetuning for downstream tasks (Dai et al., 2023). Preference tuning techniques can be used to produce further improvements (Ouyang et al., 2022; Sun et al., 2023b; Wang et al., 2024). However, most works in this area tend to propose modifications to the training algorithm and/or choice of dataset, while largely treating the pretrained components as replaceable black-box models. To complement this body of work from the efficiency point of view, we begin our investigation by analyzing the internal representations, relying on pretrained VLMs as a starting point.
Our initial inspiration comes from Huh et al. (2024), who argue that representations from different unimodal models converge towards a universal platonic representation. Although the platonic hypothesis itself is contentious, the underlying evidence suggests that model representations across modalities show increasing alignment in the deeper layers (see Figs. 4, 5 in the Appendix). This, coupled with the inner product mechanism of self-attention, invites the next question: in a multimodal setting, are cross-modal interactions more confined to deeper layers? The attention maps indeed suggest that the bulk of the activity is within the middle and latter layers (see Fig. 2). The preceding exposition culminates in our central investigation: Are the early MLLM layers necessary for multimodal token processing, and can we skip them without sacrificing performance? More specifically,
  1. 1.
    We demonstrate redundancy in early layers of multimodal LLMs, and that letting multimodal tokens bypass them entirely offers significant efficiency gains, with virtually no degradation. We refer to this framework as DeepInsert.
  2. 2.
    We showcase DeepInsert’s versatility across modalities, via popular open-source MLLMs: vision (LLaVA, BLIP (Liu et al., 2024a; Li et al., 2023a)), audio (LTU (Gong et al., 2024)), and molecules (MolCA (Liu et al., 2024e)). We also reveal a natural performance-efficiency tradeoff (see Fig. 1), enabling practitioners to select models based on their specific needs.
The remainder of the paper is organized as follows. Sec. 2 places our work in the context of broader literature. Sec. 3 motivates our framework, while Sec. 4 details the general notion of the architectural modification, highlighting implementation and design considerations. Experiments in Sec. 5 demonstrate the versatility and effectiveness across a variety of popular open-source multimodal models. Sec. 6 concludes with promising directions for broader application.
@@ 2.1 MLLMs & Alignment
The general architecture of MLLMs considered here comprises a pretrained LLM, frozen perceptual encoders for different modalities (image, audio, video, etc. (Gong et al., 2021; Radford et al., 2021; Sun et al., 2023a; Zhao et al., 2024)), and a trainable mapping module (can be a lightweight MLP or a larger Perceiver (Jaegle et al., 2021)) that aligns the modality with the LLM input space. The LLM prompts consist of text and multimodal tokens, the latter of which are obtained from the encoder via the mapping module. This general regime encompasses a wide range of works (Alayrac et al., 2022; Li et al., 2023a; Zhang et al., 2023; Bai et al., 2023; Liu et al., 2024a; Gong et al., 2024; Ghosh et al., 2024; Microsoft et al., 2025), and thus, DeepInsert is fairly universally applicable.
@@
Based on where alignment is achieved, our method relates best to the idea of model stitching (Lenc and Vedaldi, 2015; Moschella et al., 2023), that is, grafting initial layers of a neural network onto the latter half of another with an affine layer. Our work can thus be interpreted as multimodal model stitching for efficiency, without any additional bells and whistles.
2.2 Efficiency via parameter reduction
@@
Low-rank adapters (LoRA) (Hu et al., 2021; Dettmers et al., 2023) are commonly used for computationally-efficient finetuning of LLM-based models. The low-rank constraint enables finetuning large models with only a small set of additive parameters, reducing memory usage. However, inference remains just as costly. An alternative is model compression, where an existing model is effectively distilled, leading to accelerated inference and almost-at-par capabilities Fang et al. (2021); Wang et al. (2022). While these techniques exploit parametric redundancies in LLMs, our approach is motivated via functional redundancy, and therefore largely complementary.
2.3 Layer skipping & Token reduction
@@
Layer skipping, which has been employed in LLMs for speculative decoding and speeding up inference (Corro et al., 2023; Elhoushi et al., 2024; Din et al., 2024), has been recently applied to MLLMs (Shukor and Cord, 2024; Zeng et al., 2025). Specifically, Shukor and Cord (2024) combine layer skipping with compression and pruning of the base LLMs to demonstrate impressive efficiency gains in their custom setup. However, it is unclear if their method scales well as it falls $10-20\%$ short of baseline performance when trained in the LLaVA setup. On the other hand, Zeng et al. (2025) propose adaptive token skipping by selecting top-k representative tokens via cosine distances. However, their token selection is inherently LLaVA-centric, exploiting redundancies across vision tokens, and it is unclear if it generalizes to other VLMs or other modalities. FlexiDepth (Luo et al., 2025a) and AdaSkip (Luo et al., 2025b) dynamically adjust the number of layers executed based on input complexity, while MoLe-VLA (Zhang et al., 2025) and $\gamma$-MoD (Luo et al., 2024) introduce mixture-of-expert or modality-specific modules to selectively bypass computation. While effective within their respective domains, these approaches are often tailored to unimodal LLMs (FlexiDepth, AdaSkip) or require architectural modifications for multimodal settings (MoLe-VLA, $\gamma$-MoD). In contrast, DeepInsert, without relying on context-dependent routing or auxiliary modules, provides a principled framework for enabling full early layer skipping. Importantly, DeepInsert maintains strong performance across many modalities, while preserving the same training setup.
@@
Another line of research focuses on reducing the number of input tokens processed during inference such as ST3 (Zhuang et al., 2024), Skip-Vision (Zeng et al., 2025), FastV (Chen et al., 2024), VTW (Lin et al., 2024), PruMerge (Shang et al., 2024), MADTP (Cao et al., 2024), and PACT (Dhouib et al., 2025), which prune redundant tokens or learn adaptive token selection strategies. DeepInsert is conceptually orthogonal to this direction: while these are either early-exit or token pruning techniques, DeepInsert is essentially a late-entry method, which enables it to be applied in conjunction to further improve efficiency. In fact, we demonstrate in Appendix E that DeepInsert, when combined with FastV, VTW or PruMerge, either maintains or improves than baseline. This suggests that DeepInsert can make the model more robust to token pruning strategies.
@@ 3 Where Multimodal Tokens Matter
(a) Vision to language token activation in LLaVA-7B. The y-axis represents LLM depth, the x-axis consists of vision tokens arranged in order of positional indices. The score represents the relative layerwise contribution of that token.
(b) Visual Attention Ratio (VAR) (Jiang et al., 2025) in LLaVA-7B. This is the sum of attention scores from all the vision tokens to the answer token, visualized as a function of attention heads (sorted left to right in decreasing order of activity).
Figure 2: Attention activity of vision tokens, visualized pre-prediction in Fig. 2(a) and post-prediction, in Fig. 2(b), jointly indicate that vision tokens become relevant only after the initial few layers.
This section motivates our proposed framework. We first explore the nature of representational alignment in pretrained unimodal models, in the setup of Huh et al. (2024). While Figs. 4, 5 in Appendix B.1 offer evidence of emergent semantic alignment in deeper layers, it is hard to quantify its role in multimodal training. Consequently, we turn towards analyzing the self-attention maps, to shed light on how the multimodal tokens engage with the language prompt during multimodal processing. We start with a balanced subset of multiple choice questions (MCQs) drawn from the Visual7W dataset (Zhu et al., 2016), based on COCO images (Chen et al., 2015). Using MCQ-based questions allows us to first isolate data samples where the model succeeds in answering correctly by comparing the token logit against the choices (A, B, C, or D) and then, obtain the relevant attention activity for these examples, in just one forward pass. Isolating the correct instances is important, since instances of hallucination can significantly alter the attention behavior (Jiang et al., 2025).
With this setup in place, we analyze the attention activity as a function of model depth in two ways. (i) Fig. 2(a) shows the corresponding attention scores from the visual token to the last language token responsible for prediction. We subsample vision tokens (100 out of 576 tokens arranged left to right in order of positional indices) and study their layer-relative contribution, by normalizing their per-layer contribution by its net contribution across layers. This allows us to cleanly visualize tokens that contribute high absolute value of attention scores (such as CLIP summary tokens (Neo et al., 2025)) with those that do not. The attention scores correspond to the average of top-$k$ ($k=5$ heads for each layer) attention heads, with other cases in the Appendix B.2. (ii) Fig. 2(b) shows the corresponding scores with respect to the first response/answer token post prediction. Inspired by (Jiang et al., 2025), we sum up attention scores from vision tokens to the answer token across all heads. Since the total incoming attention into a given token (per head) must sum to one, we obtain a ratio of vision to language attention in each layer. If this ratio is high, it implies the answer token receives most of the attention from vision tokens, rather than language tokens. Thus this is referred to as Visual Attention Ratio (VAR) (Jiang et al., 2025). We note a consistent trend in both cases: vision tokens interact with language predominantly in intermediate layers, indicating earlier layers may be redundant in multimodal processing.
@@
Remark: In the complete layer-wise depiction (Fig. 6), there is a strong activation pattern in the first two layers. However, this is unlikely to be the key to vision-to-language information transfer. Basu et al. (2024); Jiang et al. (2025) suggest that this is LLaVA-specific and not observed for other VLMs/encoders. Moreover, our experiments in Appendix C suggest that even in LLaVA, these layers are largely redundant. So we skip the first two layers in Fig. 2 to keep the visualization cleaner.
@@ 4 Proposed Framework
These observations bring us back to our central question: Do we really need the early LLM layers to parse multimodal information? To investigate, we propose our simple yet effective modification for general MLLM frameworks, DeepInsert: inserting the multimodal tokens directly into the intermediate layers of the LLM, illustrated in Fig. 3.
4.1 Technical Challenges
Although straightforward as an abstraction, our approach poses several challenges. While multimodal tokens are inserted in the intermediate layer, language tokens must still be processed by the entire LLM for accurate language processing. Consequently, we must divide the prompt into language and multimodal constituents. The first stage of the forward pass processes just the language tokens. In the second stage, these are recombined before being processed by the latter layers of the LLM. Note that while this can be theoretically achieved by masking the vision tokens in the early layers, it naturally lowers efficiency gains and, we find, leads to noticeably worse performance. Therefore for our implementation, we need to refactor the LLM’s forward pass as well as the KV-cache structure, so the model can maintain efficient generation capabilities. For interleaved text and multimodal data, we must also ensure consistency in positional embeddings through the splitting and recombining of the prompt.
(a) Conventional architecture for MLLMs.